# Intervening nuclear obscuration changing the X-ray look of the QSO CFHQS J164121+375520

### v1

@@ 4.2 Other scenarios
In addition to intervening obscuration, V22 discussed other possibilities, such as intrinsic variability, a tidal disruption event (TDE), or gravitational lensing, as potential drivers of the variability pattern of J1641. Thanks to the new observations presented in this paper, we can now exclude intrinsic variability: simple variations of the accretion rate of J1641, either stochastic or systematic, are hardly consistent with the constant rest-frame UV emission from the QSO and, in particular, with the detection of hard-band X-ray emission with no soft-band counterpart during the recent Chandra monitoring program.
A TDE may be consistent with the bright X-ray emission in 2018 if relativistic beaming is involved (e.g., Bloom et al., 2011; Saxton et al., 2021) and the subsequent strong flux drop (e.g. Ricci et al., 2020, 2021). However, it requires a rare configuration in which the debris of the disrupted star produces heavy obscuration (e.g., Blanchard et al., 2017) to explain the detection in the hard band only in the recent Chandra dataset. Moreover, for a SMBH with the mass relevant to J1641, a TDE can only involve a giant star, for which the SMBH tidal radius is sufficiently more extended than the event horizon (e.g., MacLeod et al., 2012), thus reducing considerably the chance probability of a similar event.
If J1641 was a lensed object, as is the QSO J043947.08+163415.7 (Fan et al., 2019), the motion of the stars in the lens galaxy can produce microlensing, that affects in particular X-ray emission (e.g., Popovic et al., 2006) thus explaining the lack of strong variation of the rest-frame UV emission (e.g., Chartas et al., 2002, 2016). However, microlensing amplitudes in different X-ray bands have been found to differ by at most factors of a few (e.g., Chartas et al., 2012, 2017), which is hardly consistent with the extreme spectral variation of J1641.
@@ 5 Conclusions
J1641 is an X-ray changing look QSO that transitioned from a soft and bright state to a hard-band-only emission state, with a suppression of the soft-band flux by a factor , in few rest-frame months. While other QSOs show significant variability (Nanni et al., 2018; Moretti et al., 2021; Marcotulli et al., 2025), J1641 is by far the most extreme in terms of amplitude and spectral variation. An obscuration event happening after 2018 appears to be the most likely explanation for its X-ray and rest-frame UV lightcurves. The observed suppression of the soft-band X-ray flux requires the obscuring material to be characterized by a Compton-thick column density. Such material could be a geometrically thick inner disk itself, a dense nuclear wind, or a gas cloud orbiting the central SMBH at sub-pc/pc scales, such as the clouds forming the broad line region. We discuss how these scenarios are consistent with the lack of strong rest-frame UV variability. The population of broad-emission line AGN at high redshift recently discovered with JWST show similar properties as J1641 after the X-ray state transition, such as unobscured line emission, and weak X-ray emission (e.g., Maiolino et al., 2024), suggesting that similar physical mechanisms might be at play.
This study showcases the unique power of high-resolution and sensitive X-ray observations to investigate the nuclear properties of high-redshift QSOs, and, as a direct consequence, the physics driving the fast growth of SMBHs in the early universe. A future X-ray mission with these characteristics, such as AXIS (e.g., Marchesi et al., 2020; Reynolds et al., 2023; Cappelluti et al., 2024) is necessary to fill in the gap that Chandra will leave in the foreseeable future, and provide strong high-energy synergies to state-of-the-art and next-generation multi-wavelength facilities.
@@ Acknowledgements.
We thank the reviewer for their useful suggestions, and E. Congiu for useful discussions. We acknowledge support from the “INAF Ricerca Fondamentale” 2022 and 2023 grants. WNB acknowledges support from Chandra X-ray Center grant GO2–23080X. FEB acknowledges support from ANID-Chile BASAL CATA FB210003, FONDECYT Regular 1241005,and Millennium Science Initiative, AIM23-0001. KI acknowledges support under the grant PID2022-136827NB-C44 provided by MCIN/AEI/10.13039/501100011033 / FEDER, UE. This research has made use of data obtained from the Chandra Data Archive (Proposal ID 23700118), and software provided by the Chandra X-ray Center (CXC) in the application packages CIAO. We acknowledge the support from the LBT-Italian Coordination Facility for the execution of observations, data distribution and reduction. The LBT is an international collaboration among institutions in the United States, Italy and Germany: the University of Arizona on behalf of the Arizona university system; Istituto Nazionale di Astrofisica, Italy; LBT Beteiligungsgesellschaft, Germany, representing the Max-Planck Society, the Astrophysical Institute Potsdam, and Heidelberg University; The Ohio State University, and The Research Corporation, on behalf of The University of Notre Dame, University of Minnesota and University of Virginia. This research made use of APLpy (http://aplpy.github.com) and Astropy (Astropy Collaboration et al., 2018).References

### v2

@@ 4.2 Other scenarios
In addition to intervening obscuration, V22 discussed other possibilities, such as intrinsic variability, a tidal disruption event (TDE), or gravitational lensing, as potential drivers of the variability pattern of J1641. Thanks to the new observations presented in this Letter, we can now exclude intrinsic variability: simple variations in the accretion rate of J1641, either stochastic or systematic, are not consistent with the constant rest-frame UV emission from the QSO or, in particular, with the detection of hard-band X-ray emission with no soft-band counterpart during the recent Chandra monitoring program.
A TDE may be consistent with the bright X-ray emission in 2018, and the subsequent strong flux drop (e.g., Ricci et al., 2020, 2021), if relativistic beaming is involved (e.g., Bloom et al., 2011; Saxton et al., 2021). However, to explain the detection in the hard band only in the recent Chandra dataset, it would require a rare configuration in which the debris of the disrupted star produces heavy obscuration (e.g., Blanchard et al., 2017). Moreover, for a SMBH with a mass similar to that estimated for J1641, a TDE can only involve a giant star, for which the SMBH tidal radius is sufficiently more extended than the event horizon (e.g., MacLeod et al., 2012), thus considerably reducing the chance probability of a similar event.
If J1641 were a lensed object, as is the QSO J043947.08+163415.7 (Fan et al., 2019), the motion of the stars in the lens galaxy could produce microlensing, which would in particular affect X-ray emission (e.g., Popovic et al., 2006), thus explaining the lack of strong variation in the rest-frame UV emission (e.g., Chartas et al., 2002, 2016). However, microlensing amplitudes in different X-ray bands have been found to differ by at most factors of a few (e.g., Chartas et al., 2012, 2017), which is not consistent with the extreme spectral variation of J1641.
@@ 5 Conclusions
J1641 is an X-ray changing-look QSO that transitioned from a soft and bright state to a hard-band-only emission state, with a suppression of the soft-band flux by a factor , in few rest-frame months. While other QSOs show significant variability (Nanni et al., 2018; Moretti et al., 2021; Wolf et al., 2024; Marcotulli et al., 2025), J1641 is by far the most extreme in terms of amplitude and spectral variation. An obscuration event that occurred after 2018 appears to be the most likely explanation for its X-ray and rest-frame UV light curves. The observed suppression of the soft-band X-ray flux requires the obscuring material to be characterized by a Compton-thick column density. Such material could be a geometrically thick inner disk, a dense nuclear wind, or a gas cloud orbiting the central SMBH at parsec or subparsec scales, such as the clouds forming the broad line region. We discuss how these scenarios are consistent with the lack of strong rest-frame UV variability. The broad-emission-line AGN at high redshifts recently discovered with JWST show similar properties as J1641 after the X-ray state transition, such as unobscured line emission and weak X-ray emission (e.g., Maiolino et al., 2024), suggesting that similar physical mechanisms might be at play.
This study showcases the unique power of high-resolution and sensitive X-ray observations to investigate the nuclear properties of high-redshift QSOs and, as a direct consequence, the physics driving the fast growth of SMBHs in the early Universe. A future X-ray mission with these characteristics, such as AXIS (e.g., Marchesi et al., 2020; Reynolds et al., 2023; Cappelluti et al., 2024), is necessary to fill in the gap that Chandra will leave in the foreseeable future, and provide strong high-energy synergies to state-of-the-art and next-generation multiwavelength facilities.
@@ Acknowledgements.
We thank the reviewer for their useful suggestions, and E. Congiu for useful discussions. We acknowledge support from the “INAF Ricerca Fondamentale” 2022 and 2023 grants. WNB acknowledges support from Chandra X-ray Center grant GO2–23080X. FEB acknowledges support from ANID-Chile BASAL CATA FB210003, FONDECYT Regular 1241005,and Millennium Science Initiative, AIM23-0001. KI acknowledges support under the grant PID2022-136827NB-C44 provided by MCIN/AEI/10.13039/501100011033 / FEDER, UE. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection (CDC) “collection_id” https://doi.org/10.25574/cdc.343, and software provided by the Chandra X-ray Center (CXC) in the application packages CIAO. We acknowledge the support from the LBT-Italian Coordination Facility for the execution of observations, data distribution and reduction. The LBT is an international collaboration among institutions in the United States, Italy and Germany: the University of Arizona on behalf of the Arizona university system; Istituto Nazionale di Astrofisica, Italy; LBT Beteiligungsgesellschaft, Germany, representing the Max-Planck Society, the Astrophysical Institute Potsdam, and Heidelberg University; The Ohio State University, and The Research Corporation, on behalf of The University of Notre Dame, University of Minnesota and University of Virginia. This research made use of APLpy (http://aplpy.github.com) and Astropy (Astropy Collaboration et al., 2018).References